# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 3.1 Existence of weak solutions and conditional uniqueness
We start with proving the existence of global weak solutions to problem (3.1)–(3.3) on $[0,+\infty)$.
@@ 3.1.1 The approximate problem
Since $C^{\infty}_{0}((0,+\infty);D(\bm{S}))$ is dense in $L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))$, given $\bm{v}\in L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))$, there exists a sequence $\big{\{}\bm{v}^{k}\big{\}}\subset C^{\infty}_{0}((0,+\infty);D(\bm{S}))$ such that
@@
As a consequence, $\|\bm{v}^{k}\|_{L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))}\leq\|\bm{v}\|_{L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))}+1$ for $k$ sufficiently large. For every $k\in\mathbb{Z}^{+}$, let us consider the following approximate system with the prescribed velocity field $\bm{v}^{k}$:
@@
Since $\bm{v}^{k}$ is sufficiently regular, for arbitrary but fixed $T\in(0,+\infty)$, following the proof of [27, Lemma 3.1], we can show that problem (3.10)–(3.12) admits a unique weak solution $(\varphi^{k},\mu^{k},\sigma^{k})$ on $[0,T]$ satisfying
@@
$\varphi^{k}\in L^{\infty}(\Omega\times(0,T))\ \ \textrm{with}\ \ |\varphi^{k}|<1\quad\textrm{a.e.\ in}\ \Omega\times(0,T),$
@@
Besides, a similar argument like that in [24, Section 3] yields
$\varphi^{k}\in L^{\infty}(0,T;W^{2,6}(\Omega))\cap L^{4}(0,T;H^{2}(\Omega)),\quad\varPsi^{\prime}(\varphi^{k})\in L^{2}(0,T;L^{6}(\Omega)).$
@@ 3.1.2 Uniform estimates
We proceed by deriving estimates for the approximate solutions $(\varphi^{k},\mu^{k},\sigma^{k})$ that are uniform with respect to the parameter $k$.
@@
In a similar manner, from (3.10c) we can derive the mass conservation for $\sigma^{k}$:
@@
Second estimate. Multiplying (3.10a) by $\mu^{k}$, (3.10c) by $\sigma^{k}-\chi\varphi^{k}$, respectively, integrating over $\Omega$ and adding the resultants together, we infer from (3.10b) that
@@
$\quad+\frac{\beta}{2}\|\nabla\mathcal{N}\big{(}\varphi^{k}(t)-\overline{\varphi^{k}(t)}\big{)}\|^{2},$ (3.17)
$\mathcal{D}^{k}(t)$ $=\|\nabla\mu^{k}(t)\|^{2}+\|\nabla(\sigma^{k}(t)-\chi\varphi^{k}(t))\|^{2}.$ (3.18)
Following the argument in [30, Section 3.2], from (H2) and Hölder’s inequality, the initial energy can be estimated as follows
@@
where $C>0$ only depends on $\Omega$. On the other hand, it follows from the $L^{\infty}$-estimate (3.13) that
@@
where $C_{1}>0$ only depends on $\chi$, $\beta$ and $\Omega$.
If $\alpha=0$, we simply have $R_{1}=0$. Otherwise, the term $R_{1}$ can be estimated as in [30, (3.20)] such that
@@
where $C_{2}>0$ depends on $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$, $\Omega$ and coefficients of the system except $\alpha$. For the other two terms, using integration by parts and (3.13), we get
@@
Hence, it follows from (3.16), (3.22), (3.23) and (3.24) that
@@
for almost all $t\in[0,T]$, where
~$\widetilde{\mathcal{F}}^{k}(t)=\mathcal{F}^{k}(t)+C_{2}(1+\alpha^{3})e^{-\alpha t}|\overline{\varphi_{0}}-c_{0}|.$
Integrating in time on $[0,t]\subset[0,T]$, we obtain
~$\widetilde{\mathcal{F}}^{k}(t)+\frac{1}{4}\int_{0}^{t}\mathcal{D}^{k}(s)\,\mathrm{d}s\leq\widetilde{\mathcal{F}}^{k}(0)+(1+\chi^{2})\int_{0}^{t}\|\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s,\quad\forall\,t\in[0,T],$
@@
where $C>0$ depends on $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$, $\max_{r\in[-1,1]}|\varPsi(r)|$, $\|\varphi_{0}\|_{H^{1}}$, $\|\sigma_{0}\|$, coefficients of the system and $\Omega$, but it is independent of $k$ and $T$. Similar to the estimates [30, (3.15), (3.16)] (with $\gamma=0$ therein), we can control the total mass of the chemical potential by
$\left|\int_{\Omega}\mu^{k}\,\mathrm{d}x\right|$ $\leq C\big{(}\|\nabla\mu^{k}\|+\|\sigma^{k}-\overline{\sigma^{k}}\|+\|\varphi^{k}-\overline{\varphi^{k}}\|\big{)}\|\nabla\varphi^{k}\|+C$
@@
where $C>0$ depends on coefficients of the system, $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$ and $\Omega$. Then it follows from (3.25), (3.26) and the Poincaré–Wirtinger inequality that
$\int_{0}^{T}\|\mu^{k}(s)\|^{2}_{H^{1}}\,\mathrm{d}s\leq C\big{(}1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{4}\big{)},$ (3.27)
$\int_{0}^{T}\|\sigma^{k}(s)\|_{H^{1}}^{2}\,\mathrm{d}s\leq C\big{(}1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}\big{)},$ (3.28)
where the constant $C>0$ may depend on $T$, but still independent of $k$.
@@
It follows from [24, Lemma 7.3], (3.25) and (3.28) that
$\int_{0}^{T}\|\varphi^{k}(s)\|_{H^{2}}^{4}\,\mathrm{d}s\leq C\big{(}1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{4}\big{)}.$ (3.29)
Besides, applying [24, Lemma 7.4], we deduce from (3.25), (3.27) and (3.28) that
$\int_{0}^{T}\|\varphi^{k}(s)\|_{W^{2,6}}^{2}\,\mathrm{d}s+\int_{0}^{T}\|\varPsi^{\prime}_{0}(\varphi^{k}(s))\|_{L^{6}}^{2}\,\mathrm{d}s\leq C\big{(}1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{4}\big{)},$ (3.30)
The positive constant $C$ in (3.29) and (3.30) may depend on $T$, but it is independent of $k$.
@@
$\int_{0}^{T}\|\partial_{t}\varphi^{k}(s)\|_{(H^{1})^{\prime}}^{2}\,\mathrm{d}s$ $\leq\int_{0}^{T}\big{(}\|\bm{v}^{k}(s)\|^{2}\|\varphi^{k}(s)\|_{L^{\infty}}^{2}+\|\nabla\mu^{k}(s)\|^{2}+\alpha^{2}\|\overline{\varphi^{k}}(s)-c_{0}\|^{2}\big{)}\,\mathrm{d}s$
@@
where $C>0$ is independent of $T$ and $k$. Similarly, we can obtain the estimate for $\partial_{t}\sigma^{k}$:
$\int_{0}^{T}\|\partial_{t}\sigma^{k}(s)\|_{(H^{1})^{\prime}}^{\frac{4}{3}}\,\mathrm{d}s$
$\quad\leq\int_{0}^{T}\big{(}\|\bm{v}^{k}(s)\|_{\bm{L}^{6}}^{\frac{4}{3}}\|\sigma^{k}(s)\|_{L^{3}}^{\frac{4}{3}}+\|\nabla(\sigma^{k}-\chi\varphi^{k})(s)\|^{\frac{4}{3}}\big{)}\,\mathrm{d}s$
$\quad\leq C\sup_{t\in[0,T]}\|\sigma^{k}(t)\|^{\frac{2}{3}}\left(\int_{0}^{T}\|\bm{v}^{k}(s)\|_{\bm{H}^{1}}^{2}\,\mathrm{d}s\right)^{\frac{2}{3}}\left(\int_{0}^{T}\|\sigma^{k}(s)\|_{H^{1}}^{2}\,\mathrm{d}s\right)^{\frac{1}{3}}$
$\qquad+T^{\frac{1}{3}}\left(\int_{0}^{T}\|\nabla(\sigma^{k}-\chi\varphi^{k})(s)\|^{2}\,\mathrm{d}s\right)^{\frac{2}{3}}$
$\quad\leq C\big{(}1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}\big{)}\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{H}^{1}(\Omega))}^{\frac{4}{3}}+C\big{(}1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}\big{)},$
where $C>0$ may depend on $T$, but it is independent of $k$.
@@ 3.1.3 Existence
By the arbitrariness of $T>0$ and the uniqueness of approximate solutions $(\varphi^{k},\mu^{k},\sigma^{k})$, we can extend $(\varphi^{k},\mu^{k},\sigma^{k})$ from $[0,T]$ to $[0,+\infty)$. To construct a global weak solution to problem (3.1)–(3.3) on $[0,+\infty)$, we apply the standard compactness argument combined with a diagonal extraction process as in [8, Chapter V, Section 1.3.6] for the three dimensional Navier–Stokes equations.
In what follows, the convergence is always understood in the sense of a suitable subsequence (not relabelled for simplicity). On the time interval $[0,1]$, from the uniform estimates obtained above, we can find a triple $(\varphi_{(1)},\mu_{(1)},\sigma_{(1)})$ such that
$\varphi^{k}\to\varphi_{(1)}\quad$ weakly star $\text{weakly star in}\ L^{\infty}(0,1;H^{1}(\Omega))\cap L^{\infty}(\Omega\times(0,1)),$
$\varphi^{k}\to\varphi_{(1)}\quad$ weakly $\text{weakly in}\ L^{4}(0,1;H^{2}(\Omega))\cap L^{2}(0,1;W^{2,6}(\Omega)),$
$\partial_{t}\varphi^{k}\to\partial_{t}\varphi_{(1)}\quad$ weakly $\text{weakly in}\ L^{2}(0,1;(H^{1}(\Omega))^{\prime}),$
$\mu^{k}\to\mu_{(1)}\quad$ weakly $\text{weakly in}\ L^{2}(0,1;H^{2}(\Omega)),$
$\sigma^{k}\to\sigma_{(1)}\quad$ weakly star $\text{weakly star in}\ L^{\infty}(0,1;L^{2}(\Omega)),$
$\sigma^{k}\to\sigma_{(1)}\quad$ weakly $\text{weakly in}\ L^{2}(0,1;H^{1}(\Omega)),$
$\partial_{t}\sigma^{k}\to\partial_{t}\sigma_{(1)}\quad$ weakly $\text{weakly in}\ L^{\frac{4}{3}}(0,1;(H^{1}(\Omega))^{\prime}).$
By the Aubin–Lions compactness lemma, we also have
$\varphi^{k}\to\varphi_{(1)}\quad$ strongly $\text{strongly in}\ C([0,1];L^{2}(\Omega))\cap L^{4}(0,1;W^{1,r}(\Omega)),$
$\sigma^{k}\to\sigma_{(1)}\quad$ strongly $\text{strongly in}\ L^{2}(0,1;L^{r}(\Omega)),$
for $r\in[1,6)$, which implies
@@
As a consequence, we get
$\varPsi^{\prime}(\varphi^{k})\to\varPsi^{\prime}(\varphi_{(1)})\quad$ weakly $\text{weakly in}\ L^{2}(0,1;L^{6}(\Omega)),$
Besides, taking advantage of the strong convergence (3.9), we find that
$\bm{v}^{k}\cdot\nabla\varphi^{k}\to\bm{v}\cdot\nabla\varphi_{(1)}\quad$ weakly $\text{weakly in}\ L^{2}(0,1;L^{\frac{3}{2}}(\Omega)),$
@@
$\bm{v}^{k}\cdot\nabla\sigma^{k}\to\bm{v}\cdot\nabla\sigma_{(1)}\quad$ weakly $\text{weakly in}\ L^{1}(0,1;L^{\frac{3}{2}}(\Omega)),$
@@
Hence, we are able to pass to the limit as $k\to+\infty$ in the weak formulation for $(\varphi^{k},\mu^{k},\sigma^{k})$ and conclude that the limit triple $(\varphi_{(1)},\mu_{(1)},\sigma_{(1)})$ is a weak solution to problem (3.1)–(3.3) on $[0,1]$ with expected regularity properties. Verification of the boundary and initial conditions are straightforward.

### v2

@@ 3.1 Existence of weak solutions and conditional uniqueness
We start by proving the existence of global weak solutions to problem (3.1)–(3.3) on $0,+\infty)$.
@@ 3.1.1 The approximate problem
We first regularize the velocity field $\bm{v}\in L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))$. Since $C^{\infty}_{0}((0,+\infty);D(\bm{S}))$ is dense in $L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))$, there exists a sequence $\big\{\bm{v}^{k}\big\}\subset C^{\infty}_{0}((0,+\infty);D(\bm{S}))$ such that
@@
As a result, $\|\bm{v}^{k}\|_{L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))}\leq\|\bm{v}\|_{L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))}+1$ for $k$ sufficiently large. For every $k\in\mathbb{Z}^{+}$, consider the following approximate system with the prescribed velocity field $\bm{v}^{k}$:
@@
Since $\bm{v}^{k}$ is sufficiently regular, for arbitrary but fixed $T\in(0,+\infty)$, following the proof of [31, Lemma 3.1], we can show that problem (3.9)–(3.11) admits a unique weak solution $(\varphi^{k},\mu^{k},\sigma^{k})$ on $[0,T]$ satisfying
@@
$\varphi^{k}\in L^{\infty}(\Omega\times(0,T))\ \ \textrm{with}\ \ |\varphi^{k}(x,t)|<1\quad\textrm{a.e.\ in}\ \Omega\times(0,T),$
@@
Besides, under assumptions (H2), (H3), an argument similar to that in [27, Section 3] yields
$\varphi^{k}\in L^{2}(0,T;W^{2,6}(\Omega))\cap L^{4}(0,T;H^{2}(\Omega)),\quad\varPsi^{\prime}(\varphi^{k})\in L^{2}(0,T;L^{6}(\Omega)).$
@@ 3.1.2 Uniform estimates
Next, we derive estimates for the approximate solutions $(\varphi^{k},\mu^{k},\sigma^{k})$ that are uniform with respect to the parameter $k$ and the time. Without loss of generality, below we consider $T\geq 1$.
@@
In a similar way, from (3.9c) we can derive the mass conservation property for $\sigma^{k}$:
@@
Second estimate. Multiplying (3.9a) by $\mu^{k}$, (3.9c) by $\sigma^{k}-\chi\varphi^{k}$, respectively, integrating over $\Omega$ and adding the resultants together, from (3.9b) we infer that
@@
$\quad+\frac{\beta}{2}\|\nabla\mathcal{N}\big(\varphi^{k}(t)-\overline{\varphi^{k}(t)}\big)\|^{2},$
$\mathcal{D}^{k}(t)$ $=\|\nabla\mu^{k}(t)\|^{2}+\|\nabla(\sigma^{k}(t)-\chi\varphi^{k}(t))\|^{2}.$
Applying the argument in [34, Section 3.2], from (H2) and Hölder’s inequality, the initial energy can be estimated as follows
@@
where $C>0$ depends only on $\Omega$. On the other hand, it follows from the $L^{\infty}$-estimate (3.12) that
@@
where $C_{1}>0$ depends only on $\chi$, $\beta$, and $\Omega$.
Concerning the first term on the right-hand side of (3.15), if $\alpha=0$, we simply have $R_{1}=0$. If $\alpha>0$, $R_{1}$ can be estimated as in [34, (3.20)] such that (with $\gamma=0$ therein)
@@
where $C_{2}>0$ depends on $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$, $\Omega$ and the coefficients of the system except $\alpha$. For the other two terms, using integration by parts and (3.12), we get
@@
Hence, it follows from (3.15) and (3.19)–(3.21) that
@@
for almost all $t\in(0,T)$, where
$\widetilde{\mathcal{F}}^{k}(t)=\mathcal{F}^{k}(t)+C_{2}(1+\alpha^{3})e^{-\alpha t}|\overline{\varphi_{0}}-c_{0}|.$
Integrating in time on $0,t]\subset[0,T]$, we obtain
$\widetilde{\mathcal{F}}^{k}(t)+\frac{1}{4}\int_{0}^{t}\mathcal{D}^{k}(s)\,\mathrm{d}s\leq\widetilde{\mathcal{F}}^{k}(0)+(1+\chi^{2})\int_{0}^{t}\|\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s,\quad\forall\,t\in[0,T],$
@@
where $C>0$ depends on $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$, $\max_{r\in[-1,1]}|\varPsi(r)|$, $\|\varphi_{0}\|_{H^{1}}$, $\|\sigma_{0}\|$, the coefficients of the system, $\Omega$, but it is independent of $k$ and $T$. Similarly to the estimates [34, (3.15), (3.16)] (with $\gamma=0$ therein), we can control the integration of the chemical potential by
$\left|\int_{\Omega}\mu^{k}\,\mathrm{d}x\right|$ $\leq C\big(\|\nabla\mu^{k}\|+\|\sigma^{k}-\overline{\sigma^{k}}\|+\|\varphi^{k}-\overline{\varphi^{k}}\|\big)\|\varphi^{k}-\overline{\varphi^{k}}\|+C$
@@
where $C>0$ depends on the coefficients of the system, $\overline{\varphi_{0}}$, $\overline{\sigma_{0}}$, and $\Omega$. Then it follows from (3.22), (3.23) and the Poincaré–Wirtinger inequality that
$\sup_{t\in[0,T-1]}\int_{t}^{t+1}\|\mu^{k}(s)\|^{2}_{H^{1}}\,\mathrm{d}s\leq C\big(1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{4}\big),$ (3.24)
$\sup_{t\in[0,T-1]}\int_{t}^{t+1}\|\sigma^{k}(s)\|_{H^{1}}^{2}\,\mathrm{d}s\leq C\big(1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}\big),$ (3.25)
where the constant $C>0$ is independent of $k$, $T$ and $t$. In addition, using the interpolation $\|f\|_{L^{3}}\leq C\|f\|_{H^{1}}^{\frac{1}{2}}\|f\|^{\frac{1}{2}}$, we infer from (3.22) and (3.25) that
$\sup_{t\in[0,T-1]}\int_{t}^{t+1}\|\sigma^{k}(s)\|_{L^{3}}^{4}\,\mathrm{d}s\leq C\big(1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{4}\big),$ (3.26)
@@
It follows from [27, Lemma 7.3], (3.22) and (3.25) that
$\sup_{t\in[0,T-1]}\int_{t}^{t+1}\|\varphi^{k}(s)\|_{H^{2}}^{4}\,\mathrm{d}s\leq C\big(1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{4}\big).$ (3.27)
Moreover, applying [27, Lemma 7.4], we can deduce from (3.22), (3.24) and (3.25) that
$\sup_{t\in[0,T-1]}\int_{t}^{t+1}\|\varphi^{k}(s)\|_{W^{2,6}}^{2}\,\mathrm{d}s+\sup_{t\in[0,T-1]}\int_{t}^{t+1}\|\varPsi^{\prime}_{0}(\varphi^{k}(s))\|_{L^{6}}^{2}\,\mathrm{d}s$
$\qquad\leq C\big(1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{4}\big).$ (3.28)
The positive constant $C$ in (3.27) and (3.28) is independent of $k$, $T$ and $t$.
@@
$\int_{0}^{T}\|\partial_{t}\varphi^{k}(s)\|_{(H^{1})^{\prime}}^{2}\,\mathrm{d}s$ $\leq C\int_{0}^{T}\big(\|\bm{v}^{k}(s)\|^{2}\|\varphi^{k}(s)\|_{L^{\infty}}^{2}+\|\nabla\mu^{k}(s)\|^{2}+\alpha^{2}\|\overline{\varphi^{k}}(s)-c_{0}\|^{2}\big)\,\mathrm{d}s$
@@
where $C>0$ is independent of $k$ and $T$. Similarly, using (3.9c), (3.22) and (3.26), we can obtain the following estimate for $\partial_{t}\sigma^{k}$:
$\int_{t}^{t+1}\|\partial_{t}\sigma^{k}(s)\|_{(H^{1})^{\prime}}^{\frac{4}{3}}\,\mathrm{d}s$
$\quad\leq C\int_{t}^{t+1}\big(\|\bm{v}^{k}(s)\|_{\bm{L}^{6}}^{\frac{4}{3}}\|\sigma^{k}(s)\|_{L^{3}}^{\frac{4}{3}}+\|\nabla(\sigma^{k}-\chi\varphi^{k})(s)\|^{\frac{4}{3}}\big)\,\mathrm{d}s$
$\quad\leq C\left(\int_{0}^{T}\|\bm{v}^{k}(s)\|_{\bm{H}^{1}}^{2}\,\mathrm{d}s\right)^{\frac{2}{3}}\left(\int_{t}^{t+1}\|\sigma^{k}(s)\|_{L^{3}}^{4}\,\mathrm{d}s\right)^{\frac{1}{3}}+\left(\int_{t}^{t+1}\|\nabla(\sigma^{k}-\chi\varphi^{k})(s)\|^{2}\,\mathrm{d}s\right)^{\frac{2}{3}}$
$\quad\leq C\big(1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}\big)\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{H}^{1}(\Omega))}^{\frac{4}{3}}+C\big(1+\|\bm{v}^{k}\|_{L^{2}(0,T;\bm{L}^{2}(\Omega))}^{2}\big),$
for all $t\in[0,T-1]$, where $C>0$ is independent of $k$, $T$ and $t$.
@@ 3.1.3 Existence
For any $k\in\mathbb{Z}^{+}$, due to the arbitrariness of $T>0$ and the uniqueness of the approximate solution $(\varphi^{k},\mu^{k},\sigma^{k})$, we can extend $(\varphi^{k},\mu^{k},\sigma^{k})$ from $[0,T]$ to the entire interval $0,+\infty)$. To construct a global weak solution to the original problem (3.1)–(3.3) on $0,+\infty)$, we apply the standard compactness argument combined with a diagonal extraction process as in [10, Chapter V, Section 1.3.6] for the three-dimensional Navier–Stokes equations.
In what follows, the convergence is always understood in the sense of a suitable subsequence (not relabeled for simplicity). On the time interval $[0,1]$, from the uniform estimates obtained above, we can find a triple $(\varphi_{(1)},\mu_{(1)},\sigma_{(1)})$ such that
$\varphi^{k}\to\varphi_{(1)}\quad$ $\text{weakly star in}\ L^{\infty}(0,1;H^{1}(\Omega))\cap L^{\infty}(\Omega\times(0,1)),$
$\varphi^{k}\to\varphi_{(1)}\quad$ $\text{weakly in}\ L^{4}(0,1;H^{2}_{N}(\Omega))\cap L^{2}(0,1;W^{2,6}(\Omega)),$
$\partial_{t}\varphi^{k}\to\partial_{t}\varphi_{(1)}\quad$ $\text{weakly in}\ L^{2}(0,1;(H^{1}(\Omega))^{\prime}),$
$\mu^{k}\to\mu_{(1)}\quad$ $\text{weakly in}\ L^{2}(0,1;H^{1}(\Omega)),$
$\sigma^{k}\to\sigma_{(1)}\quad$ $\text{weakly star in}\ L^{\infty}(0,1;L^{2}(\Omega)),$
$\sigma^{k}\to\sigma_{(1)}\quad$ $\text{weakly in}\ L^{2}(0,1;H^{1}(\Omega))\cap L^{4}(0,1;L^{3}(\Omega)),$
$\partial_{t}\sigma^{k}\to\partial_{t}\sigma_{(1)}\quad$ $\text{weakly in}\ L^{\frac{4}{3}}(0,1;(H^{1}(\Omega))^{\prime}).$
By the Aubin–Lions compactness lemma, we have
$\varphi^{k}\to\varphi_{(1)}\quad$ $\text{strongly in}\ L^{\infty}(0,1;H^{1-r}(\Omega))\cap L^{4}(0,1;H^{2-r}(\Omega)),$
$\sigma^{k}\to\sigma_{(1)}\quad$ $\text{strongly in}\ L^{2}(0,1;H^{1-r}(\Omega)),$
for any $r\in(0,1/2)$, which further implies
@@
The convergence of singular nonlinear terms $\big\{\varPsi^{\prime}(\varphi^{k})\big\}$ can be achieved using the argument in [3, Section 2]. We sketch it here for the convenience of the readers. From the pointwise convergence of $\varphi^{k}$ and the fact that $|\varphi^{k}(x,t)|<1$ almost everywhere in $\Omega\times(0,1)$, we find $\varphi_{(1)}\in L^{\infty}(\Omega\times(0,1))$ with $|\varphi_{(1)}(x,t)|\leq 1$ almost everywhere in $\Omega\times(0,1)$. Due to (H2), i.e., the continuity of $\varPsi^{\prime}$ in (1,1)(-1,1), we see that $\varPsi^{\prime}(\varphi^{k})\to\widetilde{\varPsi}^{\prime}(\varphi_{(1)})$ almost everywhere in $\Omega\times(0,1)$, where $\widetilde{\varPsi}^{\prime}(s)=\varPsi^{\prime}(s)$ if $s\in(-1,1)$ and $\widetilde{\varPsi}^{\prime}(\pm 1)=\pm\infty$. Hence, an application of Fatou’s lemma yields
$\int_{\Omega\times(0,1)}|\widetilde{\varPsi}^{\prime}(\varphi_{(1)})|^{2}\,\mathrm{d}x\mathrm{d}t\leq\liminf_{k\to+\infty}\int_{\Omega\times(0,1)}|\varPsi^{\prime}(\varphi^{k})|^{2}\,\mathrm{d}x\mathrm{d}t<+\infty,$
so that $\widetilde{\varPsi}^{\prime}(\varphi_{(1)})\in L^{2}(\Omega\times(0,1))$, which implies $|\varphi_{(1)}(x,t)|<1$ almost everywhere in $\Omega\times(0,1)$. As a result, $\widetilde{\varPsi}^{\prime}(\varphi_{(1)})=\varPsi^{\prime}(\varphi_{(1)})$ holds almost everywhere in $\Omega\times(0,1)$. This observation, together with the uniform boundedness of $\big\{\varPsi^{\prime}(\varphi^{k})\big\}$ in $L^{2}(0,1;L^{6}(\Omega))$ implies that
$\varPsi^{\prime}(\varphi^{k})\to\varPsi^{\prime}(\varphi_{(1)})\quad\text{weakly in}\ L^{2}(0,1;L^{6}(\Omega)).$
Next, from the strong convergence (3.8) for $\bm{v}^{k}$, the weak/strong convergence properties of $\sigma^{k}$ and the incompressibility condition for $\bm{v}^{k}$, $\bm{v}$, we find
$\left|\int_{0}^{1}\!\!\int_{\Omega}(\bm{v}^{k}\cdot\nabla\sigma^{k}-\bm{v}\cdot\nabla\sigma_{(1)})\xi\,\mathrm{d}x\mathrm{d}t\right|$
@@
$\quad\leq\left|\int_{0}^{1}\!\!\int_{\Omega}\sigma^{k}(\bm{v}^{k}-\bm{v})\cdot\nabla\xi\,\mathrm{d}x\mathrm{d}t\right|+\left|\int_{0}^{1}\!\!\int_{\Omega}(\sigma^{k}-\sigma_{(1)})\bm{v}\cdot\nabla\xi\,\mathrm{d}x\mathrm{d}t\right|$
$\quad\leq\|\sigma^{k}\|_{L^{\infty}(0,1;L^{2}(\Omega))}^{\frac{1}{2}}\|\sigma^{k}\|_{L^{2}(0,1;H^{1}(\Omega))}^{\frac{1}{2}}\|\bm{v}^{k}-\bm{v}\|_{L^{2}(0,1;\bm{L}^{6}(\Omega))}\|\nabla\xi\|_{L^{4}(0,1;\bm{L}^{2}(\Omega))}$
$\qquad+\left|\int_{0}^{1}\!\!\int_{\Omega}(\sigma^{k}-\sigma_{(1)})\bm{v}\cdot\nabla\xi\,\mathrm{d}x\mathrm{d}t\right|$
$\quad\to 0,\quad\text{as}\ k\to+\infty,$
for any $\xi\in L^{4}(0,1;H^{1}(\Omega))$. Here, for the convergence of the second integral we have used the weak convergence $\sigma^{k}\to\sigma_{(1)}$ in $L^{4}(0,1;L^{3}(\Omega))$ and the fact $\bm{v}\cdot\nabla\xi\in L^{\frac{4}{3}}(0,1;L^{\frac{3}{2}}(\Omega))$. Similarly, one can check that
$\left|\int_{0}^{1}\!\!\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{k}-\bm{v}\cdot\nabla\varphi_{(1)})\zeta\,\mathrm{d}x\mathrm{d}t\right|\to 0,\quad\text{as}\ k\to+\infty,$
@@
for any $\zeta\in L^{2}(0,1;H^{1}(\Omega))$. Hence, we are able to pass the limit as $k\to+\infty$ in the weak formulation for $(\varphi^{k},\mu^{k},\sigma^{k})$ and conclude that the limit triple $(\varphi_{(1)},\mu_{(1)},\sigma_{(1)})$ is a weak solution to problem (3.1)–(3.3) on the time interval $[0,1]$ with expected regularity properties. Verification of the initial conditions is straightforward (cf. also Remark 3.1).